# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 2. PRELIMINARIES
To see that the converse of the above implication fails, consider strongly inaccessible cardinals $\kappa$ and $\lambda$ with $\kappa<\lambda$ and $\kappa=$ the first strongly inaccessible cardinals. Then $\left(\mathrm{V}_{\kappa},\in\right)$ satisfies the $\Pi_{1}$-statement expressing “there is no strongly inaccessible cardinal”, but $\left(\mathrm{V}_{\lambda},\in\right)$ satisfies the negation of the same statement.
(j) As we shall see in Lemma 4.3 faithful end extensions of models of $\mathrm{ZF}$ are rank extensions.
(k) Thanks to the provability of the induction scheme (over natural numbers) in $\mathrm{ZF}$, if $\mathcal{N}$ is a faithful extension of a model $\mathcal{N}$ (where $\mathcal{M}$ is a model of $\mathrm{ZF}$), then $\mathcal{N}$ fixes each element of $\omega^{\mathcal{M}}$ (but $\mathcal{N}$ need not fix $\omega^{\mathcal{M}}$, e.g., consider the case when $\mathcal{N}$ is an internal ultrapower of $\mathcal{M}$ modulo a nonprincipal ultrafilter over $\omega^{\mathcal{M}}$).
2.7. Definition. Reasoning within $ZF$, for each object $a$ in the universe of sets, let $c_{a\text{ }}$ be a constant symbol denoting $a$ (where the map $a\mapsto c_{a}$ is $\Delta_{1}).$ For each finite extension $\mathcal{L}$ of $\mathcal{L}_{\mathrm{Set}}$, let $\mathrm{Sent}_{\mathcal{L}^{+}}(x)$ be the $\mathcal{L}_{\mathrm{Set}}$-formula that defines the class $\mathrm{Sent}_{\mathcal{L}^{+}}$ of sentences in the language $\mathcal{L}^{+}=$$\mathcal{L}\cup\{c_{a}:a\in\mathrm{V}\}$, and let $\mathrm{Sent}_{\mathcal{L}^{+}}(i,x)$ be the $\mathcal{L}_{\mathrm{set}}$-formula that expresses “$i\in\omega,$ $x\in\mathrm{Sent}_{\mathcal{L}^{+}},$ and $x$ is a $\Sigma_{i}(\mathcal{L}^{+})$-sentence”. Suppose $\mathcal{L}$ is a finite extension of $\mathcal{L}_{\mathrm{Set}}$, $\mathcal{M}\models\mathrm{ZF}\mathcal{(L)}$, $S\subseteq M$, and $k\in\omega^{\mathcal{M}}$.
@@
$(i)\ \ \left[\left(S\left(c_{x}=c_{y}\right)\leftrightarrow x=y\right)\wedge\left(S\left(c_{x}\in c_{y}\right)\leftrightarrow x\in y\right)\right].$
$(ii)\ \ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\lnot\psi\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\lnot S\mathsf{(}\psi\mathsf{)}\right]\mathsf{.}$
$(iii)$ $\ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\psi_{1}\vee\psi_{2}\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\left(S(\psi_{1})\vee S(\psi_{2})\right)\right]\mathsf{.}$
$(iv)$ $\ \left[\mathrm{Sent}_{\mathcal{L}^{+}}(k,\varphi)\wedge\left(\varphi=\exists v\ \psi(v)\right)\right]\rightarrow\left[S(\varphi)\leftrightarrow\exists x\ S\mathsf{(}\psi\mathsf{(}c_{x}\mathsf{))}\right].$
(b) $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$ if for each $k\in\omega,$ $S$ is a $\Sigma_{k}$-satisfaction class over $\mathcal{M}$.
(a) (Definability of Truth) If $\mathcal{M}$ is coded as an element of a model $\mathcal{N}$ of a sufficiently strong [^8] The two canonical fragments of ZF that are capable of defining Tarskian truth predicates for their internal set structures are KP (see [Bar, III.1]; the proof makes it clear that only $\Pi_{1}$-Foundation is invoked) and the fragment of Zermelo set theory in which the separation scheme is limited to $\Delta_{0}$-formulae(see [M, Proposition 3.10]). fragment of $\mathrm{ZF}$, then for some $\mathcal{N}$-coded $S$, $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$ (we refer to $S$ as the Tarskian satisfaction class for $\mathcal{M}$, as defined in $\mathcal{N}$).
(b) (Undefinability of Truth) If $S$ is a $\Sigma_{\omega}$-satisfaction class for $\mathcal{M}$, then $S$ is not $\mathcal{M}$-definable
It is a well-known result of Levy that if $\mathcal{M}\models\mathrm{ZF,}$ then there is a $\Delta_{0}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ both by a $\Sigma_{1}$-formula and a $\Pi_{1}$-formula (see [J, p. 186] for a  proof). This makes it clear that for each $n\geq 1,$ there is a $\Sigma_{n}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ by a $\Sigma_{n}$-formula. Levy’s result extends to models of $\mathrm{ZF(}\mathcal{L}\mathrm{)}$ if $\mathcal{L}$ is finite as  follows:
[…]
The Reflection Theorem is often formulated as a theorem scheme of $\mathrm{ZF}$ (e.g., as in [J]), the proof strategy of the Reflection Theorem applies equally well to $\mathrm{ZF}(\mathcal{L})$ for finite extensions $\mathcal{L}$ of $\mathcal{L}_{\mathrm{set}}$.
2.10. Montague-Vaught-Levy Reflection Theorem. Let $\mathcal{L}$ be a finite extension of $\mathcal{L}_{\mathrm{set}}.$ For each $n\in\omega$, $\mathrm{ZF(}\mathcal{L}\mathrm{)}$ proves that there are arbitrarily large ordinals $\gamma$ such that the submodel of the universe determined by $V$γ is a $\Sigma_{n}(\mathcal{L})$-elementary submodel of the universe. [^9] This is expressible with the help of $\mathrm{Sat}_{\Sigma_{n}(\mathcal{L})}.$
@@
The following is a special case of a result of Gaifman [G-1, Theorem 1, p.54].
@@
2.12. Gaifman Splitting Theorem. Let $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{Set}}.$ Suppose $\mathcal{M}\models\mathrm{ZF}(\mathcal{L})$, $\mathcal{N}$ is an $\mathcal{L}$-structure with $\mathcal{M}\preceq_{\Delta_{0}(\mathcal{L})}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Then the following hold:
(a) $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{{}_{\Sigma_{0}(\mathcal{L}),\mathrm{end}}}\mathcal{N}$.
(b) If $\mathcal{M\preceq N}$, then $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{{}_{\mathrm{end}}}\mathcal{N}$.
Proof. To simplify the notation, we present the proof for $\mathcal{L}=\mathcal{L}_{\mathrm{set}}$ since the same reasoning handles the general case. To prove (a) suppose $\mathcal{M}\prec_{\Delta_{0},\mathrm{cof}}\mathcal{N},$ where $\mathcal{M}\models\mathrm{ZF}\mathsf{.}$ Note that we are only assuming that $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-structure such that $\mathcal{M}$ is a submodel of $\mathcal{N}$, and we are not assuming that $\mathcal{N}\models\mathrm{ZF}\mathsf{.}$ The definition of $\mathcal{M}^{\ast}$ makes it clear that $\mathcal{M}$ is cofinal in $\mathcal{N}$. Our goal is to show that $\mathcal{M}\prec_{\Sigma_{n}}\mathcal{M}^{\ast}$ for each $n\in\omega,$ so let us fix $n\in\omega.$ Thanks to the Reflection Theorem there is an unbounded subset $U$ of $\mathrm{Ord}^{\mathcal{M}}$ such that:
(1)   If $\alpha$ and $\beta$ are in $U$ with $\alpha<\beta$, then $\mathcal{M}\models\mathrm{V}_{\alpha}\prec_{\Sigma_{n}}\mathrm{V}_{\beta}.$
Let $\mathrm{Sat}(s,y)$ be the $\mathcal{L}_{\mathrm{set}}$-formula expressing “$s$ is the Tarskian satisfaction predicate for $(y,\in)$”. $\mathrm{Sat}(s,y)$ is $\Delta_{0}.$ Since in $\mathrm{ZF}$ for each $y$ there is a (unique) $s$ such that $\mathrm{Sat}(s,y)$ holds, this makes it clear that the statement expressing $(x,\in)\prec_{\Sigma_{n}}(y,\in)$ can be written as a $\Delta_{0}$-statement in parameters $x$, $y$, and $s$. Thus,
(2)   If $m_{1}$ and $m_{2}$ are elements of $M$ such that $\mathcal{M}\models(m_{1},\in)\prec_{\Sigma_{n}}(m_{2},\in)$, and $\mathcal{M}\prec_{\Delta_{0}}\mathcal{M}^{\ast}$, then $\mathcal{M}^{\ast}\models(m_{1},\in)\prec_{\Sigma_{n}}(m_{2},\in).$
Let $\alpha\in\mathrm{Ord}^{\mathcal{M}},$ let $a_{\alpha}\in M$ such that $\mathcal{M}\models a_{\alpha}=\mathrm{V}_{\alpha}$,  using (1) and (2) we can conclude:
@@
(3)   If $\alpha$ and $\beta$ are in $U$ with $\alpha<\beta$, then $\left(a_{\alpha},\in\right)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\left(a_{\beta},\in\right)^{\mathcal{M}^{\ast}}.$
On the other hand, by putting $\mathcal{M}\prec_{\mathrm{cof}}\mathcal{M}^{\ast}$ together with $\mathcal{M}=\bigcup\limits_{\alpha\in U}(a_{\alpha},\in)^{\mathcal{M}}$ we have:
(4) $\ \ \mathcal{M}^{\ast}=\bigcup\limits_{\alpha\in U}(a_{\alpha},\in)^{\mathcal{M}^{\ast}}$.
Thanks to (3), (4), and the Elementary Chains Theorem we conclude:
(5)   If $\alpha$ is in $U$, then $(a_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\mathcal{M}^{\ast}.$
Thus, given a $\Sigma_{n}$-formula $\varphi(\overline{x})$ such that $\mathcal{M}\models\varphi(\overline{p})$ for some sequence $\overline{p}$ of parameters from $M$, there is a large enough $\alpha\in U$ such that:
$\mathcal{M}\models\overline{p}\in a_{\alpha}\wedge\varphi^{a_{\alpha}}(\overline{p}).$
Since the statement $\varphi^{a_{\alpha}}(\overline{p})$ is $\Delta_{0}$, $\mathcal{M}^{\ast}\models\varphi^{a_{\alpha}}(\overline{p})$, i.e., $(a_{\alpha},\in)^{\mathcal{M}^{\ast}}\models\varphi(\overline{p})$, which together with (5) implies that $\mathcal{M}^{\ast}\models\varphi(\overline{a}).$ This concludes the proof of (a).
(b) follows easily from (a).
$\square$
@@
2.13. Remark. In general, if $\mathcal{M}\subseteq\mathcal{N}$, where $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathrm{ZF}$, the condition “$\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$” does not imply that $\mathcal{M}$ is cofinal in $\mathcal{N}$ (for example consider $\mathcal{M}$ and $\mathcal{N}$ where $\mathcal{N}$ is a set-generic extension of $\mathcal{M}$). However, the conjunction of $\Sigma_{1}$-elementarity and “$\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$”, where $\mathcal{M}\models\mathrm{ZF}$ (or even $\mathrm{KPR}$) and $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-set structure implies that $\mathcal{M}$ is cofinal in $\mathcal{N}$, since the statement $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ is $\Pi_{1}$, and therefore it is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}.$ [^10] It is well-known that for models $\mathcal{M}$ and $\mathcal{N}$ of $\mathrm{ZF}$, $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}$ iff $\mathcal{M}\prec_{\Delta_{0}^{\mathcal{P}}}\mathcal{N}$. For a refinement of the nontrivial direction of this result, see [M, Proposition 6.17]. This is how the existence of elementary embedding between inner models of $\mathrm{ZF}$ is usually formalized, as in Kanamori’s monograph [Kan, Proposition 5.1].
@@
Recall that for models of $\mathrm{ZF}$, the $\mathcal{L}_{\mathrm{set}}$-sentence $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ expresses: “there is some $p$ such that every set is first order definable in some structure of the form $\left(\mathrm{V}_{\alpha},\in,p,\beta\right)_{\beta<\alpha}$ with $p\in\mathrm{V}_{\alpha}$”. The following result is classical.
@@
2.14. Myhill-Scott Theorem. The following statements are equivalent for $\mathcal{M}\models\mathrm{ZF}$:
(a) $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$

### v6

@@ 2. PRELIMINARIES
  1. (a)
    Given $k\in\omega$, $S$ is a $\Sigma_{k}$-satisfaction class for $\mathcal{M}$ if $\left(\mathcal{M},S\right)\models\mathrm{Sat}(k,\mathrm{S})$, where $\mathrm{Sat}(k,\mathrm{S})$ is the universal generalization of the conjunction of the axioms $(i)$ through $(iv)$ below. We assume that first order logic is formulated using only the logical constants $\left\{\lnot,\vee,\exists\right\}.$
    1. $(i)$
      $\left[\left(\mathrm{S}\left(\dot{x}=\dot{y}\right)\leftrightarrow x=y\right)\wedge\left(\mathrm{S}\left(\dot{x}\in\dot{y}\right)\leftrightarrow x\in y\right)\right].$
    2. $(ii)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\lnot\psi\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\lnot\mathrm{S}\mathsf{(}\psi\mathsf{)}\right]\mathsf{.}$
    3. $(iii)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\psi_{1}\vee\psi_{2}\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\left(\mathrm{S}(\psi_{1})\vee\mathrm{S}(\psi_{2})\right)\right]\mathsf{.}$
    4. $(iv)$
      $\left[\mathrm{Sent}_{\mathcal{L}_{\mathrm{set}}}(k,\varphi)\wedge\left(\varphi=\exists v\ \psi(v)\right)\right]\rightarrow\left[\mathrm{S}(\varphi)\leftrightarrow\exists x\ \mathrm{S}\mathsf{(}\psi\mathsf{(}\dot{x}\mathsf{))}\right].$
  2. (b)
    $S$ is a full satisfaction class for $\mathcal{M}$ if $\left(\mathcal{M},S\right)\models\mathrm{Sat}(k,\mathrm{S})$ for every $k\in\omega$.
  1. (a)
    (Definability/Codability of Truth) If $\mathcal{M}$ is a structure coded as an element $m$ of a model $\mathcal{N}$ of a sufficiently strong [^12] There are two canonical fragments of $\mathsf{ZF}$ that are ‘sufficiently strong’ for this purpose, namely: (1) $\mathsf{KP}$ (see Definition 2.1(d)), as shown by Friedman, Lu, and Wong in [FLW, Lemma 4.1]. (2) The fragment $\mathsf{M}_{0}$ of $\mathsf{Z}$ (see Definition 2.1(e)), as shown by Mathias [M-1, Proposition 3.10]. Note, however, that much weaker systems suffice if one only wishes to have a set theory within which the Tarskian satisfaction relation of every internal set structure is definable, as opposed to: coded as a set. One such weak system is $\mathsf{DS}$ (for ‘Devlin strengthened’), which is shown by Mathias [M-2, Proposition 10.37] to be capable of defining the Tarskian satisfaction predicate for set structures. fragment of $\mathsf{ZF}$, then there is $s\in N$ such that $\mathcal{N}$ views $s$ as the Tarskian satisfaction relation on $m$, written $\mathcal{N}\models\mathrm{sat}(s,m).$ [^13] In other words, within $\mathcal{N}$, $s$ is the set consisting of ordered pairs $(\varphi(\vec{x}),\vec{a})$ such that $m\models\varphi\left(\vec{x}/\vec{a}\right).$ Moreover, $\mathrm{sat}(x,y)$ is $\Delta_{1}$ in $\mathcal{N}.$
  2. (b)
    (Undefinability of Truth) If $S$ is a full satisfaction class for $\mathcal{M}$, then $S$ is not $\mathcal{M}$-definable.
It is a well-known result of Levy that if $\mathcal{M}\models\mathsf{ZF}\mathrm{,}$ then there is a $\Delta_{0}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ both by a $\Sigma_{1}$-formula and a $\Pi_{1}$-formula (see [J, p. 186] for a proof). This makes it clear that for each $n\geq 1,$ there is a $\Sigma_{n}$-satisfaction class for $\mathcal{M}$ that is definable in $\mathcal{M}$ by a $\Sigma_{n}$-formula. Levy’s result extends to models of $\mathsf{ZF}\mathrm{(}\mathcal{L}\mathrm{)}$ if $\mathcal{L}$ is finite as follows:
@@ 4. $(iv)$
The Reflection Theorem is often formulated as a theorem scheme of $\mathsf{ZF}$ (e.g., as in [J]), the proof strategy of the Reflection Theorem applies equally well to $\mathsf{ZF}(\mathcal{L})$ for finite extensions $\mathcal{L}$ of $\mathcal{L}_{\mathrm{set}}$.
2.10. Montague-Vaught-Levy Reflection Theorem. Let $\mathcal{L}$ be a finite extension of $\mathcal{L}_{\mathrm{set}}.$ For each $n\in\omega$, $\mathsf{ZF}(\mathcal{L})$ proves that there are arbitrarily large ordinals $\gamma$ such that the submodel of the universe determined by VVγ is a $\Sigma_{n}(\mathcal{L})$-elementary submodel of the universe. [^15] This is expressible with the help of $\mathrm{Sat}_{\Sigma_{n}(\mathcal{L})}.$
@@
The following is a special case of a result of Gaifman [G-1, Theorem 1, p.54]. In the statement and the proof, $\mathcal{M}^{\ast}\prec_{{}_{\Delta_{0}(\mathcal{L}),\mathrm{end}}}\mathcal{N}$ is shorthand for the conjunction of $\mathcal{M}^{\ast}\prec_{\Delta_{0}(\mathcal{L)}}\mathcal{N}$ and $\mathcal{M}^{\ast}\prec_{\mathrm{end}}\mathcal{N}$.
@@
2.12. Gaifman Splitting Theorem. Let $\mathcal{L}\supseteq\mathcal{L}_{\mathrm{set}}.$ Suppose $\mathcal{M}\models\mathsf{ZF}(\mathcal{L})$, $\mathcal{N}$ is an $\mathcal{L}$-structure with $\mathcal{M}\prec_{\Delta_{0}(\mathcal{L})}\mathcal{N}$, and $\mathcal{M}^{\ast}$ is the convex hull of $\mathcal{M}$ in $\mathcal{N}$. Then the following hold:
  1. (a)
    $\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{\Delta_{0}(\mathcal{L}),\mathrm{end}}\mathcal{N}$.
  2. (b)
    $\mathcal{M\prec N}\Rightarrow\mathcal{M}\preceq_{\mathrm{cof}}\mathcal{M}^{\ast}\preceq_{\mathrm{end}}\mathcal{N}$.
Proof. To simplify the notation, we present the proof for $\mathcal{L}=\mathcal{L}_{\mathrm{set}}$ since the same reasoning handles the general case. In our proof we use the notation $\varphi^{v}$, where $\varphi$ is an $\mathcal{L}_{\mathrm{set}}$-formula and $v$ is a parameter, to refer to the $\Delta_{0}$-formula obtained by restricting all the quantifiers of $\varphi$ to the elements of $v$. A straightforward induction on the complexity of formulae shows that:
(1) If $\mathcal{K}$ is an $\mathcal{L}_{\mathrm{set}}$-structure, and $m\in K$, then $\mathcal{K}\models\varphi^{m}$ iff $(m,\in)^{\mathcal{K}}\models\varphi$.
Note that in the above, $\varphi$ is allowed to contain parameters in $\mathrm{Ext}_{\mathcal{K}}(m).$ Suppose $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N},$ where $\mathcal{M}\models\mathsf{ZF.}$ Note that we are not assuming that $\mathcal{N}\models\mathsf{ZF}$. Putting (1) together with $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N}$ implies that $(m,\in)^{\mathcal{M}}\preceq(m,\in)^{\mathcal{M}^{\ast}}$ for all $m\in M.$ On the other hand, the definition of $\mathcal{M}^{\ast}$ makes it clear that $(m,\in)^{\mathcal{N}}=(m,\in)^{\mathcal{M}^{\ast}}$ if $m\in M.$ Putting all this together, we have:
(2) For all $m\in M$, $(m,\in)^{\mathcal{M}}\preceq(m,\in)^{\mathcal{M}^{\ast}}=(m,\in)^{\mathcal{N}}$.
It is easy to see, using the definition of $\mathcal{M}^{\ast}$, that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}\subseteq_{\mathrm{end}}\mathcal{N}$, and in particular $\mathcal{M}^{\ast}\preceq_{\Delta_{0}}\mathcal{N}$ by Remark 2.6(b). Thus the proof of (a) is complete once we show that $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}$ for each $n\in\omega.$ Towards this aim, given $n\in\omega$, we can invoke the Reflection Theorem in $\mathcal{M}$ to get hold of some $U\subseteq\mathrm{Ord}^{\mathcal{M}}$ such that:
@@
(3) $(\mathcal{M},U)$ satisfies “$U$ is unbounded in $\mathrm{Ord},$ and $\left[\left(\mathrm{V}_{\alpha},\in\right)\prec_{\Sigma_{n}}\left(\mathrm{V},\in\right)\right]$ for every $\alpha\in U$.
For each $\alpha\in U$ let $m_{\alpha}\in M$ such that $m_{\alpha}=\mathrm{V}_{\alpha}^{\mathcal{M}}.$ By (3), we have:
(4) If $\alpha,\beta\in U$ with $\alpha<\beta$, then $(m_{\alpha},\in)^{\mathcal{M}}\prec_{\Sigma_{n}}(m_{\beta},\in)^{\mathcal{M}}.$
Next, we observe that if $a,b$ are in $M$ with $(a,\in)^{\mathcal{M}}\prec_{\Sigma_{n}}(b,\in)^{\mathcal{M}}$, then for each $k$-ary $\Sigma_{n}$-formula $\sigma(\vec{x})$, $\mathcal{M}$ satisfies the $\Delta_{0}$-statement $\delta_{\sigma}(a,b),$ where:
$\delta_{\sigma}(a,b):=\forall\vec{x}\in a\ \left[\sigma^{a}(\vec{x})\leftrightarrow\sigma^{b}(\vec{x})\right]$,
where the prefix $\forall\vec{x}\in a$ is shorthand for $\forall x_{0}\in a\cdot\cdot\cdot\forall x_{k-1}\in a.$ Therefore, by putting (4) and (2) together with the assumption $\mathcal{M}\prec_{\Delta_{0}}\mathcal{N}$ we can conclude:
(5) If $\alpha,\beta\in U$ with $\alpha<\beta$, then $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}(m_{\beta},\in)^{\mathcal{M}^{\ast}}.$
On the other hand, since $U$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}$, the definition of $m_{\alpha}$, together with the fact that $\mathcal{M}\subseteq_{\mathrm{cof}}\mathcal{M}^{\ast}$, imply:
[…]
Thanks to (2), (5), (6), and Theorem 2.11 ($\Sigma_{n}$-Elementary Chains Theorem) we can conclude:
(7) If $\alpha\in U$, then $(m_{\alpha},\in)^{\mathcal{M}}\preceq(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}\mathrm{.}$
We can conclude that $\mathcal{M}\preceq_{\Sigma_{n}}\mathcal{M}^{\ast}$ by (4), (7), and Theorem 2.11. This concludes the proof of (a).
To prove (b), suppose $\mathcal{M}\prec\mathcal{N}.$ With (a) at hand, we only need to show that $\mathcal{M}^{\ast}\preceq_{\Sigma_{n}}\mathcal{N}$ for each $n\in\omega$. We observe that if $\alpha\in U,$ then by (5) $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}\prec_{\Sigma_{n}}\mathcal{M}^{\ast},$ and by (2), $(m_{\alpha},\in)^{\mathcal{M}^{\ast}}=(m_{\alpha},\in)^{\mathcal{N}}.$ On the other hand, the assumption $\mathcal{M}\prec\mathcal{N}$ together with (3) makes it clear that if $\alpha\in U,$ $(m_{\alpha},\in)^{\mathcal{N}}\prec_{\Sigma_{n}}\mathcal{N}$. Thus $\mathcal{M}^{\ast}\preceq_{\Sigma_{n}}\mathcal{N}$ by Theorem 2.11. $\square$
@@
2.13. Remark. In general, if $\mathcal{M}\subseteq\mathcal{N}$, where $\mathcal{M}$ and $\mathcal{N}$ are both models of $\mathsf{ZF}$, the condition $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$” does not imply that $\mathcal{M}$ is cofinal in $\mathcal{N}$ (for example consider $\mathcal{M}$ and $\mathcal{N}$ where $\mathcal{N}$ is a set-generic extension of $\mathcal{M}$). However, the conjunction of $\Sigma_{1}$-elementarity and $\mathrm{Ord}^{\mathcal{M}}$ is cofinal in $\mathrm{Ord}^{\mathcal{N}}$”, where $\mathcal{M}\models\mathsf{ZF}$ (or even $\mathsf{KPR}$) and $\mathcal{N}$ is an $\mathcal{L}_{\mathrm{set}}$-structure implies that $\mathcal{M}$ is cofinal in $\mathcal{N}$, since the statement $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ can be written as a $\Pi_{1}$-statement within $\mathsf{KPR}$, and therefore it is absolute between $\mathcal{M}$ and $\mathcal{N}$ if $\mathcal{M}\prec_{\Sigma_{1}}\mathcal{N}.$ [^16] As noted in [Kan, Lemma 0.2, p5], the rank function is $\Delta_{1}^{\mathsf{ZF}}$, and this fact can be used to show that $(y\in\mathrm{Ord}\wedge x=\mathrm{V}_{y})$ can be written as a $\Pi_{1}$-statement within $\mathsf{ZF}$; an examination of the proof makes it clear that $\mathsf{KPR}$ suffices for this purpose. This is how elementary embeddings between inner models are usually formalized within $\mathsf{ZF}$, as in Kanamori’s monograph [Kan, Proposition 5.1].
@@
Recall that for models of $\mathsf{ZF}$, the $\mathcal{L}_{\mathrm{set}}$-sentence $\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ expresses: “there is some $p$ such that every set is first order definable in some structure of the form $\left(\mathrm{V}_{\alpha},\in,p,\beta\right)_{\beta<\alpha}$ with $p\in\mathrm{V}_{\alpha}$”. The following result is classical.
@@
2.14. Myhill-Scott Theorem [MS]The following statements are equivalent for $\mathcal{M}\models\mathsf{ZF}$:
  1. (a)
    $\mathcal{M}\models\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right).$
  2. (b)